# Linear and nonlinear instability of vortex columns

### v2

@@ 2.2. The outer equation
where $\widetilde{\chi}_{\rm out}:=1-\eta_{\widetilde{\ell}_{\rm out}}$, and ~$\widetilde{\ell}_{\rm out}:=\ell_{\rm out}/2$, so that $\widetilde{\chi_{\rm out}}=1$ on $\operatorname{supp}\,\chi_{\rm out}$. Thus we can consider, instead of the second equation in (2.4), the problem
@@
Thanks to Lemma 2.9, we will be able to find a linear combination (2.58) of the real and imaginary parts of
@@
If $|r-r_{0}|\in[Cn^{-3/4},C^{-1}n^{-1/2}]$, then $-ib_{0}^{1/2}$ is the dominant term in the above expansion, and so, using the expansion $(1+x)^{-2}=1-2x+O(x^{2})$ for $|x|<1$, we obtain
$\frac{1}{\gamma^{2}}=\frac{1}{b_{0}}\left(-1+i\frac{\Lambda^{\prime\prime}(r_{0})}{b_{0}^{1/2}}n(r-r_{0})^{2}+nO(|r-r_{0}|^{3}+n^{-3/2}+n|r-r_{0}|^{4})\right)$
@@
We conclude this section with verification that the solution $\upsilon$ to (2.55) depends holomorphically on $\widehat{\omega}$. To this end, we extract from $\widetilde{k}$ the main part $\widetilde{k_{\mathrm{app}}}$, which is independent of $\widehat{\omega}$; namely, we define $\widetilde{k_{\mathrm{app}}}$ by (2.54), except that $\omega$ (in the definition (1.18) of $k$) is replaced by $\omega-\widehat{\omega}=n\Lambda_{0}+ib_{0}^{1/2}+\mu_{m}$ (recall (1.27)). We can thus rewrite (2.55) as
$-\upsilon^{\prime\prime}+\widetilde{k_{\mathrm{app}}}\upsilon=f+\left(\widetilde{k_{\mathrm{app}}}-\widetilde{k}\right)\upsilon.$ (2.63)
@@
where $C^{-1}$ can be made arbitrarily small provided that $\widehat{\omega}\ll n^{-1/2}$ and $n$ is sufficiently large. Define the multiplication operator $K\colon Z\to L^{2}$, $K\upsilon:=\left(\widetilde{k_{\mathrm{app}}}-\widetilde{k}\right)\upsilon$, and $S\colon L^{2}\to Z$ the solution operator to $-\upsilon^{\prime\prime}+\widetilde{k_{\mathrm{app}}}\upsilon=g$ (i.e., $Sg=\upsilon$). Then we see that any solution $\upsilon$ of (2.55) satisfies
@@
i.e., $\upsilon=(I-SK)^{-1}Sf$, where the invertibility of $I-SK$ follows from the facts that $\|S\|_{L^{2}\to N}\lesssim n^{-3/4}$ by (2.62) and $n^{-3/4}\|K\|_{Z\to L^{2}}\leq C^{-1}$. This verifies that the solution in Lemma 2.10 depends holomorphically on $\widehat{\omega}\in D(0,\varepsilon n^{-1/2})$ for sufficiently small $\varepsilon$ and sufficiently large $n$.
@@ 2.3. The projected system
$\begin{cases}&\Psi^{\prime\prime}-(K_{0}+K_{2}\xi^{2})\Psi=Q\left(\widetilde{X}_{\rm in}K_{\rm err}(W+\Psi)-R\phi_{\rm out}X_{\rm out}^{\prime\prime}-2n^{-3/4}R\phi_{\rm out}^{\prime}X_{\rm out}^{\prime}\right)\\ &\int\Psi\overline{W}=0\\ &\phi_{\rm out}^{\prime\prime}-\widetilde{k}\phi_{\rm out}=-2\phi_{\rm in}^{\prime}\chi_{\rm in}^{\prime}-\phi_{\rm in}\chi_{\rm in}^{\prime\prime}\,.\\ \end{cases}$ (2.65)
@@
There exist $D_{\rm out},\widehat{D}\geq 1$ and $D_{\rm in}\geq 2D_{\rm out}$ such that for all $\widehat{\omega}_{m}\in B(\widehat{D}^{-1}n^{-1/2})$ and $n$ sufficiently large, the projected system (2.65) is uniquely solvable in the class $(\Psi,\phi_{\rm out})\in Y_{w}\times Z$ with the choices of cut-off scales
@@
$\|\phi_{\rm out}\|_{Z}\lesssim(\ell_{\rm in}n^{3/4})^{-1}\exp\left(-C^{-1}(\ell_{\rm in}n^{3/4})^{2}\right)\,.\hskip 176.407pt$ (2.68)
@@
$\begin{split}\mathsf{A}\colon Y_{w}\to Y_{w}\,,&\quad\mathsf{A}F=G\Leftrightarrow G^{\prime\prime}-(K_{0}+K_{2}\xi^{2})G=Q(\widetilde{X}_{\rm in}K_{\rm err}F)\,,\;\int G\overline{W}=0\\ \mathsf{B}\colon Z\to Y_{w}\,,&\quad\mathsf{B}f=G\Leftrightarrow G^{\prime\prime}-(K_{0}+K_{2}\xi^{2})G=Q(-RfX_{\rm out}^{\prime\prime}-2n^{-3/4}Rf^{\prime}{X}_{\rm out}^{\prime\prime})\,,\;\int G\overline{W}=0\\ \mathsf{C}\colon Y_{w}\to Z\,,&\quad\mathsf{C}F=g\Leftrightarrow g^{\prime\prime}-\widetilde{k}g=-2n^{3/4}R^{-1}F^{\prime}\chi_{\rm in}^{\prime}-R^{-1}F\chi_{\rm in}^{\prime\prime}\,.\end{split}$ (2.70)
@@
The parameter $D_{\rm in}$ enforces a separation between the inner scale $n^{-3/4}$ and the inner cut-off scale $\ell_{\rm in}$. The operator $\mathsf{C}$ benefits from the Gaussian weight when there is a separation of scales, since it only sees the ‘tails’ of the inner function. However, the inner equation in (2.65) depends on the error $K_{\rm err}$ of the approximation of $K$ by $K_{0}+K_{2}\xi^{2}$, and so our control of $K_{\rm err}$ and, hence, the operator $\mathsf{A}$ become worse when $\ell_{\rm in}$ is increased. Since $\mathsf{A}W$ and $\mathsf{C}W$ enter the right-hand side of (2.71), the different choices of $\ell_{\rm in}$ produce a range of estimates on the solutions. With uniqueness, the solutions to (1.17) will satisfy all the estimates simultaneously, with the worse of the two estimates holding in the overlap region. [^8] The Gaussian weight is not technically necessary to prove existence in this problem, but it yields more precise estimates.
@@
$K_{\rm err}=in^{1/2}{\widehat{\omega}}\frac{2p_{0}}{{b_{0}}^{1/2}}+(1+|\xi|^{4})O(n^{-1/2})\,,$ (2.75)
for $|\widehat{\omega}|\leq n^{-1/2}$ and $|\xi|\leq n^{1/4}$, see (A.3) for justification.
@@
$\begin{split}\|G\|_{Y_{w}}&\lesssim\|Q(-RfX_{\rm out}^{\prime\prime}-n^{-3/4}Rf^{\prime}X_{\rm out}^{\prime})\|_{L^{2}_{w}}\\ &\lesssim n^{3/8}\left(\|f\|(\ell_{\rm out}n^{3/4})^{-2}+n^{-3/4}\|f^{\prime}\|(\ell_{\rm out}n^{3/4})^{-1}\right)\\ &\lesssim\|f\|_{Z}\,,\end{split}$ (2.78)
@@
and applying (2.73) and (2.74). ∎
The operators $\mathsf{A}$ and $\mathsf{C}$ depend analytically on $\widehat{\omega}$ (through $K_{\rm err}$ and $\widetilde{k}$, respectively), and therefore the solutions to the projected system (2.65) also depend analytically on $\widehat{\omega}$.
@@ 2.4. The one-dimensional reduced equation
In order to complete the existence part of Theorem 1.1, it remains to solve the reduced equation (2.19), that is, to find $\widehat{\omega}=\widehat{\omega}(n)$ such that
@@
$|h|$ $=n^{-1/2}|\int_{\mathbb{R}}\left(\kappa\widetilde{X}_{\rm in}(W+\Psi)-R\phi_{\rm out}X_{\rm out}^{\prime\prime}-2n^{-3/4}R\phi_{\rm out}^{\prime}X_{\rm out}^{\prime}\right)W|$ (2.89)
@@ 2.5. Uniqueness
We now demonstrate the uniqueness part of Theorem 1.1, for which we use a gluing procedure that is somewhat dual to the gluing procedure for existence: [^10] The cutting operator $\phi\mapsto(\phi\chi_{\rm in},\phi\chi_{\rm out})$ in the uniqueness proof is the $L^{2}$ adjoint of the operator $(\phi_{\rm in},\phi_{\rm out})\mapsto\phi_{\rm in}\chi_{\rm in}+\phi_{\rm out}\chi_{\rm out}$, which sums the components into the ansatz for the existence proof. For linear gluing problems involving self-adjoint operators, one expects that the gluing systems for existence and uniqueness will be formal adjoints. For our purposes, it is enough to note that the roles of the cut-offs $\chi_{\rm in}$, $\chi_{\rm out}$ appearing in the inner-outer system (2.4) are switched. Suppose that $\phi\in H^{1}_{0}(\mathbb{R}^{+})$ is a solution to (1.17) with $\omega=n\Lambda(r_{0})+ib_{0}^{1/2}+\mu$ for some $\mu\in B(C_{M}n^{-1/2})$. We decompose
@@
$\begin{cases}&\phi_{\rm in}^{\prime\prime}-k\phi_{\rm in}=2\phi_{\rm out}^{\prime}\chi_{\rm in}^{\prime}+\phi_{\rm out}\chi_{\rm in}^{\prime\prime}\text{on }{\rm supp}\,\chi_{\rm in}\\ &\phi_{\rm out}^{\prime\prime}-k\phi_{\rm out}=2\phi_{\rm in}^{\prime}\chi_{\rm out}^{\prime}+\phi_{\rm in}\chi_{\rm out}^{\prime\prime}\text{ on }{\rm supp}\,\chi_{\rm out}\,,\end{cases}$ (2.93)
nearly identical to (2.4); notably, the cutoffs $\chi_{\rm in}$ and $\chi_{\rm out}$ in the boundary terms are switched. Uniqueness for solutions to (2.93), up to multiplication by constants, will imply uniqueness for solutions to (1.17).
@@
The proof is divided into two cases. When the eigenvalue is close to an approximate eigenvalue, we proceed as before, applying the Lyapunov-Schmidt reduction and uniquely solving the equations, so that $\phi$ is a constant multiple of a constructed solution. When the eigenvalue is not close to one of the approximate eigenvalues, the inner operator will be invertible without projecting the equation, and we demonstrate that $\phi=0$.
@@
$2\ell_{\rm out}=\ell_{\rm in}\,,\quad\ell_{\rm out}=Dn^{-3/4}$ (2.94)
for sufficiently large $D$.
@@
We proceed as in the proof of existence. For brevity, we will omit writing the index $m$. We let $k_{0}$, $k_{2}$ be defined as in (1.22) with $\mu$ replaced by $\mu_{m}$. Let $\Psi=\Phi_{\rm in}-P\Phi_{\rm in}$, where $PF=(\int F\overline{W})W$ denotes the $L^{2}$-orthogonal projection onto the Weber function $W$, so that $P\Psi=0$. After multiplying $\phi$ by a suitable constant, we may assume that $\Phi_{\rm in}=\iota W+\Psi$ with $\iota\in\{0,1\}$. Suppose $\iota=1$. (When $\iota=0$ it is not even necessary to consider the one-dimensional reduced equation.) We define analogues of the operators in (2.70) by
$\begin{split}\widetilde{\mathsf{A}}\colon Y\to Y,&\quad\widetilde{\mathsf{A}}F=G\Leftrightarrow G^{\prime\prime}-(K_{0}+K_{2}\xi^{2})G=Q(\widetilde{X}_{\rm in}K_{\rm err}F)\,,\;\int G\overline{W}=0\\ \widetilde{\mathsf{B}}\colon Z\to Y,&\quad\widetilde{\mathsf{B}}f=G\Leftrightarrow G^{\prime\prime}-(K_{0}+K_{2}\xi^{2})G=Q(RfX_{\rm in}^{\prime\prime}+2n^{-3/4}Rf^{\prime}{X}_{\rm in}^{\prime\prime})\,,\;\int G\overline{W}=0\\ \widetilde{\mathsf{C}}\colon Y\to Z,&\quad\widetilde{\mathsf{C}}F=g\Leftrightarrow g^{\prime\prime}-\widetilde{k}g=2n^{3/4}R^{-1}F^{\prime}\chi_{\rm out}^{\prime}+R^{-1}F\chi_{\rm out}^{\prime\prime}\,.\end{split}$ (2.95)
@@
[…]
@@
We next analyze the reduced equation (2.97). Recall the analogy with (2.87). The main point is to estimate the analytic function
@@
where in the second term we exploited the cut-offs: $|X^{\prime}_{\rm out}|,|X^{\prime\prime}_{\rm out}|\lesssim D^{-1}$. As before, we may choose $R_{0}=\widehat{D}^{-1}n^{-1/2}$ and close Rouché’s theorem in the disc $D(R_{0})$. Therefore, there is a unique choice of $\widehat{\omega}$ and $(\Psi,\phi_{\rm out})$, as desired.
@@
Nonlinear instability arguments often begin with linear semigroup estimates. For vortex columns, the main difficulty is that the abstract spectral theory only guarantees $L^{Q}$ semigroup estimates Fourier mode-by-Fourier mode, not uniformly in the modes. We circumvent this difficulty by (i) developing high-order approximations which decay backward-in-time at any prescribed rate, and (ii) controlling the growth of the remainder by naïve methods (energy estimates). This method is due to Grenier [25]. The key point is that the construction of the approximate solution only requires the linear semigroup estimates for finitely many modes depending on the desired order of the approximation.

### v3

@@ 2.2. The outer equation
where $\widetilde{\chi}_{\rm out}:=1-\eta_{\widetilde{\ell}_{\rm out}}$, and $\widetilde{\ell}_{\rm out}:=\ell_{\rm out}/2$, so that $\widetilde{\chi_{\rm out}}=1$ on $\operatorname{supp}\,\chi_{\rm out}$. Thus we can consider, instead of the second equation in (2.4), the problem
@@
Thanks to Lemma 2.9 below, we will be able to find a linear combination (2.59) of the real and imaginary parts of
@@
If $|r-r_{0}|\in[Cn^{-3/4},C^{-1}n^{-1/2}]$, then $ib_{0}^{1/2}$ is the dominant term in the above expansion, and so, using the expansion $(1+x)^{-2}=1-2x+O(x^{2})$ for $|x|<1$, we obtain
$\frac{1}{\gamma^{2}}=\frac{1}{b_{0}}\left(-1+i\frac{\Lambda^{\prime\prime}(r_{0})}{b_{0}^{1/2}}n(r-r_{0})^{2}+nO(|r-r_{0}|^{3}+n^{-3/2}+n|r-r_{0}|^{4})\right),$
@@
We conclude this section with a verification that the solution $\upsilon$ to (2.56) is holomorphic with respect to $\widehat{\omega}$. To this end, we extract from $\widetilde{k}$ the main part $\widetilde{k}_{\mathrm{app}}$, which is independent of $\widehat{\omega}$; namely, we define $\widetilde{k}_{\mathrm{app}}$ by (2.55), except that $\omega$ (in the definition (1.19) of $k$) is replaced by $\omega-\widehat{\omega}=n\Lambda_{0}+ib_{0}^{1/2}+\mu_{m}$ (recall (1.28)). We can thus rewrite (2.56) as
$\upsilon^{\prime\prime}+\widetilde{k}_{\mathrm{app}}\upsilon=f+\left(\widetilde{k}_{\mathrm{app}}-\widetilde{k}\right)\upsilon.$ (2.64)
@@
where $C^{-1}$ can be made arbitrarily small provided that $\widehat{\omega}\ll n^{-1/2}$ and $n$ is sufficiently large. Define the multiplication operator $K\colon Z\to L^{2}$, $K\upsilon:=\left(\widetilde{k}_{\mathrm{app}}-\widetilde{k}\right)\upsilon$, and $S\colon L^{2}\to Z$ the solution operator to $\upsilon^{\prime\prime}+\widetilde{k}_{\mathrm{app}}\upsilon=g$ (i.e., $Sg=\upsilon$). Then we see that any solution $\upsilon$ of (2.56) satisfies
@@
i.e., $\upsilon=(I-SK)^{-1}Sf$, where the invertibility of ISKI-SK follows from the facts that $\|S\|_{L^{2}\to N}\lesssim n^{-3/4}$ by (2.63) and $n^{-3/4}\|K\|_{Z\to L^{2}}\leq C^{-1}$. This verifies that the solution in Lemma 2.10 is holomorphic with respect to $\widehat{\omega}\in D(0,\varepsilon n^{-1/2})$ for sufficiently small $\varepsilon$ and sufficiently large $n$.
@@ 2.3. The projected system
[…]
@@
There exist $D_{\rm out},\widehat{D}\geq 1$ and $D_{\rm in}\geq 2D_{\rm out}$ such that for all $\widehat{\omega}\in B(\widehat{D}^{-1}n^{-1/2})$ and $n$ sufficiently large, the projected system (2.66) is uniquely solvable in the class $(\Psi,\phi_{\rm out})\in Y_{w}\times Z$ with the choices of cut-off scales
@@
$\|\phi_{\rm out}\|_{Z}\lesssim(\ell_{\rm in}n^{3/4})^{-1}\exp\left(-C^{-1}(\ell_{\rm in}n^{3/4})^{2}\right)\hskip 176.407pt$ (2.69)
and is holomorphic with respect to $\widehat{\omega}$ in the above norms.
@@
$\begin{split}\mathsf{A}\colon Y_{w}\to Y_{w}\,,&\quad\mathsf{A}F=G\Leftrightarrow G^{\prime\prime}-(K_{0}+K_{2}\xi^{2})G=Q(\widetilde{X}_{\rm in}K_{\rm err}F)\,,\;\int GW^{*}=0\\ \mathsf{B}\colon Z\to Y_{w}\,,&\quad\mathsf{B}f=G\Leftrightarrow G^{\prime\prime}-(K_{0}+K_{2}\xi^{2})G=Q(-RfX_{\rm out}^{\prime\prime}-2n^{-3/4}Rf^{\prime}{X}_{\rm out}^{\prime\prime})\,,\;\int GW^{*}=0\\ \mathsf{C}\colon Y_{w}\to Z\,,&\quad\mathsf{C}F=g\Leftrightarrow g^{\prime\prime}-\widetilde{k}g=-2n^{3/4}R^{-1}F^{\prime}\chi_{\rm in}^{\prime}-R^{-1}F\chi_{\rm in}^{\prime\prime}\,.\end{split}$ (2.71)
@@
The parameter $D_{\rm in}$ enforces a separation between the inner scale $n^{-3/4}$ and the inner cut-off scale $\ell_{\rm in}$. The operator $\mathsf{C}$ benefits from the Gaussian weight when there is a separation of scales, since it only sees the ‘tails’ of the inner function. However, the inner equation in (2.66) depends on the error $K_{\rm err}$ of the approximation of $K$ by $K_{0}+K_{2}\xi^{2}$, and so our control of $K_{\rm err}$, as well as the operator $\mathsf{A}$, become worse when $\ell_{\rm in}$ is increased. Since $\mathsf{A}W$ and $\mathsf{C}W$ enter the right-hand side of (2.72), the different choices of $\ell_{\rm in}$ produce a range of estimates on the solutions. With uniqueness, the solutions to (1.18) will satisfy all the estimates simultaneously, with the worse of the two estimates holding in the overlap region. [^8] The Gaussian weight is not technically necessary to prove existence in this problem, but it yields more precise estimates.
@@
$K_{\rm err}=in^{1/2}{\widehat{\omega}}\frac{2p_{0}}{{b_{0}}^{1/2}}+(1+|\xi|^{4})O(n^{-1/2})\qquad\text{ as }n\to\infty\,$, (2.76)
for $|\widehat{\omega}|\leq n^{-1/2}$ and $|\xi|\leq n^{1/4}$, see (A.3) for a justification.
@@
[…]
@@
and applying (2.74) and (2.75). The operators $\mathsf{A}$ and $\mathsf{C}$ are holomorphic with respect to $\widehat{\omega}$ (through $K_{\rm err}$ and $\widetilde{k}$, respectively), and therefore the solutions to the projected system (2.66), obtained via the Neumann series, are also holomorphic with respect to $\widehat{\omega}$. ∎
@@ 2.4. The one-dimensional reduced equation
In order to complete the proof of existence part of Theorem 1.1, it remains to solve the reduced equation (2.19), that is, to find $\widehat{\omega}=\widehat{\omega}(n)$ such that
@@
$|h|$ $=n^{-1/2}\left|\int_{\mathbb{R}}\left(\kappa\widetilde{X}_{\rm in}(W+\Psi)-R\phi_{\rm out}X_{\rm out}^{\prime\prime}-2n^{-3/4}R\phi_{\rm out}^{\prime}X_{\rm out}^{\prime}\right)W\right|$ (2.90)
@@ 2.5. Uniqueness
We now demonstrate the uniqueness part of Theorem 1.1, for which we use a gluing procedure that is somewhat dual to the gluing procedure for existence: [^10] The cutting operator $\phi\mapsto(\phi\chi_{\rm in},\phi\chi_{\rm out})$ in the uniqueness proof is the $L^{2}$ adjoint of the operator $(\phi_{\rm in},\phi_{\rm out})\mapsto\phi_{\rm in}\chi_{\rm in}+\phi_{\rm out}\chi_{\rm out}$, which sums the components into the ansatz for the existence proof. For linear gluing problems involving self-adjoint operators, one expects that the gluing systems for existence and uniqueness will be formal adjoints. For our purposes, it is enough to note that the roles of the cut-offs $\chi_{\rm in}$, $\chi_{\rm out}$ appearing in the inner-outer system (2.4) are swapped. Suppose that $\phi\in H^{1}_{0}(\mathbb{R}^{+})$ is a solution to (1.18) with $\omega=n\Lambda(r_{0})+ib_{0}^{1/2}+\mu$ for some $\mu\in B(C_{M}n^{-1/2})$. We decompose
@@
[…]
nearly identical to (2.4); notably, the cutoffs $\chi_{\rm in}$ and $\chi_{\rm out}$ in the boundary terms are swapped. Uniqueness for solutions to (2.94), up to multiplication by constants, will imply uniqueness for solutions to (1.18).
@@
The proof is divided into two cases. When the eigenvalue is close to an approximate eigenvalue, we proceed as in the existence part, applying the Lyapunov-Schmidt reduction and uniquely solving the equations, so that $\phi$ is a constant multiple of a constructed solution. When the eigenvalue is not close to one of the approximate eigenvalues, the inner operator will be invertible without projecting the equation, and we demonstrate that $\phi=0$.
@@
$\ell_{\rm in}:=2\ell_{\rm out}\,,\quad\ell_{\rm out}:=Dn^{-3/4}$ (2.95)
for sufficiently large $D>0$.
@@
We proceed as in the proof of existence. For brevity, we will omit writing the index $m$. We let $k_{0}$, $k_{2}$ be defined as in (1.23) with $\mu$ replaced by $\mu_{m}$. Let $\Psi=\Phi_{\rm in}-P\Phi_{\rm in}$, where $PF=(\int FW^{*})W$ denotes the $L^{2}$-orthogonal projection onto the Weber function $W$, so that $P\Psi=0$. After multiplying $\phi$ by a suitable constant, we may assume that $\Phi_{\rm in}=\iota W+\Psi$ with $\iota\in\{0,1\}$. Suppose $\iota=1$. (When $\iota=0$ it is not even necessary to consider the one-dimensional reduced equation.) We define analogues of the operators in (2.71) by
$\begin{split}\widetilde{\mathsf{A}}\colon Y\to Y,&\quad\widetilde{\mathsf{A}}F=G\Leftrightarrow G^{\prime\prime}-(K_{0}+K_{2}\xi^{2})G=Q(\widetilde{X}_{\rm in}K_{\rm err}F)\,,\;\int GW^{*}=0\\ \widetilde{\mathsf{B}}\colon Z\to Y,&\quad\widetilde{\mathsf{B}}f=G\Leftrightarrow G^{\prime\prime}-(K_{0}+K_{2}\xi^{2})G=Q(RfX_{\rm in}^{\prime\prime}+2n^{-3/4}Rf^{\prime}{X}_{\rm in}^{\prime\prime})\,,\;\int GW^{*}=0\\ \widetilde{\mathsf{C}}\colon Y\to Z,&\quad\widetilde{\mathsf{C}}F=g\Leftrightarrow g^{\prime\prime}-\widetilde{k}g=2n^{3/4}R^{-1}F^{\prime}\chi_{\rm out}^{\prime}+R^{-1}F\chi_{\rm out}^{\prime\prime}\,.\end{split}$ (2.96)
@@
We define the auxiliary space $\mathsf{N}:=Y\times Z$ and norm $\|(F,f)\|_{\mathsf{N}}:=\|F\|_{Y}+E\|f\|_{Z}$, as in (2.73) with $Y$ replacing $Y_{w}$. The parameter $E$ is chosen so that $E^{-1}\|\widetilde{\mathsf{B}}\|_{Z\to Y}\leq 1/6$. [^11] Alternatively, unlike in the existence proof, one actually gets $D^{-1}$ in this term. Then the restriction $D\gg 1$ is applied so that $E\|\widetilde{\mathsf{C}}\|_{Y\to Z}\leq 1/6$. The restrictions that $n\gg 1$ (depending on $D$) and $\widehat{D}\gg 1$ (independently of $D$), guarantee that $\|\widetilde{\mathsf{A}}\|_{Y\to Y}\leq 1/6$. Hence, since the operator norm $\|\mathsf{T}\|_{\mathsf{N}\to\mathsf{N}}\leq 1/2$, the projected system (2.97) is uniquely solvable, the solution is holomorphic with respect to $\widehat{\omega}$, and
@@
We next analyze the reduced equation (2.98). Recall the analogy with (2.88). The main point is to estimate the holomorphic function
@@
where in the second term we exploited the cut-offs: $|X^{\prime}_{\rm out}|,|X^{\prime\prime}_{\rm out}|\lesssim D^{-1}$. As before, we may choose $R_{0}=\widehat{D}^{-1}n^{-1/2}$ and close Rouché’s theorem in the disc $D(R_{0})$. Therefore, there is a unique choice of $\widehat{\omega}$ and $(\Psi,\phi_{\rm out})$. A straightforward computation shows that the solution constructed in Sections 2.12.4 also solves (2.108), and so it must equal to $\phi$, as desired.
@@
Nonlinear instability arguments often begin with linear semigroup estimates. For vortex columns, the main difficulty is that the abstract spectral theory only guarantees $L^{Q}$ semigroup estimates Fourier mode-by-Fourier mode, not uniformly in the modes. We circumvent this difficulty by (i) developing high-order approximations which decay backward-in-time at any prescribed rate, and (ii) controlling the growth of the remainder by naïve methods (energy estimates). This method is due to Grenier [25]. The key point is that the construction of the approximate solution only requires the linear semigroup estimates for finitely many iterations depending on the desired order of the approximation.